Scenario #1 200 Scenario #2 Scenario #3 Suppose a price-discriminating monopoly has segregated its market into two sub-markets (Market 1 and Market 2) and can prevent resale between the two. Assume that its marginal cost is $10 and equal to its average total cost of 10. The firm's demand schedule for the first group is glven by the first two columns of the table Market 1 Market 2 18 14 72 518 10 12 70 18 $90 72 70 12 72 I$12)
5. In Scenario #1, is the monopoly operating at a profit or loss? 6. In Scenario # 1, what is the monopoly's operating total proft or loss amount in dollars? Write out how you computed this amount. 7. What is the profit maximizing output (quantity) for Scenario #2? 8. what is the price will be charged in Scenario #2? 9. In Scenario #2, is the monopoly operating at a total profit or loss or zero profit? Write out how you computed this amount. 10. In Scenario #2, what is the monopoly's operating profit or loss amount in dollars? Write out how you computed this amount. 11. In Scenario #3, what output level and price will maximize the frm's profit in Market 1? 12. In Scenario #3, what output level and price will maximize the firm's profit in Market 2? 13. Based on your answers for question 11 & 12, which market is more elastic demand? 14. What is this firm's total economic profit if it price discriminates?
Answer #1

A monopoly maximizes profit by producing at the point where marginal revenue equals marginal cost. Monopoly Sets its profit maximizing price at the point where profit maximizing quantity lies on the demand curve.

5. In scenario 1, the monopoly's profit maximizing quantity is 6 units and profit maximizing price is $100. At the profit maximizing quantity, ATC = $75. As at the profit maximizing quantity, price > ATC, it means the monoply is operating at a profit.

6. Total revenue = price * quantity = $(100*6) = $600 and total cost = ATC * quantity = $(75*6) = $450. Therefore, profit = TR - TC = $(600 - 450) = $150.

7. In scenario 2, profit maximizing quantity is 4 units, l

8. In scenario 2, profit maximizing quantity lies on the demand curve at Price = $100. So, profit maximizing price is $100.

9. At the profit maximizing quantity, ATC = $125 which is greater than profit maximizing price. It means Total cost will exceed total revenue at profit maximizing quantity. Therefore, the monoply is operating at a loss = $(125*4) - $(100*4) = $100.

10. In scenario 2, the monopoly's operating loss = $100. (Calculated in previous part).